# A Weighted Loss Approach to Robust Federated Learning under Data Heterogeneity

### v2

@@ 6.2 Theoretical Results
7 Experimental Setup
This section describes our experimental setup, including the baseline methods, the WoLA approach under training objective attack, and the datasets used. We detail the configuration of the heterogeneity regimes, workers, and the types of Byzantine attacks we consider. We also present the robust aggregation methods, evaluation metrics, optimization algorithms, models, and hyperparameters we used. Finally, we explain how we ensure the reproducibility of our experiments.
7.1 Baseline Approaches
We compare WoLA against three baselines: Bucketing (BKT) [24], NNM [3], and FoundationFL (FoundFL) [15]. BKT requires setting a bucket size $s$. Following the authors’ recommendation, we use $s=\left\lfloor{n}/({2f})\right\rfloor$. When $f>n/4$, the bucket size drops to $1$, making BKT ineffective. For instance, with $17$ workers, BKT has no effect when $f=6$ or $8$. FoundFL requires a hyperparameter $m$. As suggested in the original paper, we set $m=n/2$, i.e., $m=8$ in all our experiments. We also include a baseline using only a robust aggregator without any additional defense, denoted as -.
We evaluate different variants of WoLA. In particular, to assess robustness against training objective attacks, we implement the worst-case scenario described in Subsection 7.2, denoting the resulting loss as WoLA. Additionally, we examine the combination of WoLA and WoLA with NNM.
We do not compare WoLA with methods that (i) require a server-side dataset [4, 28, 11, 10, 30], (ii) involve high computational costs [4, 12, 32, 47, 39], or (iii) rely on hyperparameter tuning [46, 35].
7.2 WoLA under Training Objective Attack
To define the training objective, we consider a scenario where workers share their local label distributions with the server. In this case, Byzantine machines may try to manipulate the objective by sending malicious distributions. We analyze the impact of such attacks.
Just as Byzantine workers can send arbitrary updates with full knowledge of honest gradients, we assume they can also send arbitrary distributions with full knowledge of local label distributions. Rather than listing all possible attacks, we focus on one that causes the maximum distortion: Byzantine machines send a zero vector except for a single entry set to $1$, corresponding to the minority class in the global dataset (proof are provided in Appendix A.7).
To defend against attacks, we propose using the geometric median as the aggregator. This approach offers two key advantages: (i) Unlike methods like Trimmed Mean or Krum, which exclude outliers, the geometric median treats all updates equally. This is important because outlier distributions are not necessarily malicious. Excluding them may discard honest distribution and degrade performance. (ii) When aggregating probability distributions, the geometric median produces a probability distribution—unlike the coordinate-wise median, for instance—ensuring the training objective retains its expected form.
Our goal is to show that a strong attack, even when countered by a simple defense, has limited impact on model performance. As the chosen attack represents a worst-case scenario, it provides a meaningful benchmark for evaluating training objective attacks.
We refer to our approach in this scenario as WoLA.
7.3 Datasets, Heterogeneity & Workers
We conduct the experiments on four datasets: MNIST [27], Fashion-MNIST [42], CIFAR10 [26], and Purchase100 [37]. Purchase100 contains shopping records for thousands of customers. We use a simplified version [37], where each customer is represented by $600$ binary features indicating item purchases. The data is clustered into $100$ classes, and the task is to predict each customer’s class.
To simulate data heterogeneity across workers, we sample from the global dataset using a Dirichlet distribution with concentration parameter $\alpha$, following the method in [3, 23]. We consider three heterogeneity levels—low, medium, and high—corresponding to $\alpha\in\{3,1,0.3\}$.
We use $n=17$ workers in all experiments. The number of Byzantine workers varies as $f\in\{2,4,6,8\}$. Each worker holds a local dataset of fixed size, equal to the total dataset size divided by the number of honest workers.
7.4 Byzantine Attacks & Robust Aggregators
Byzantine workers carry out one of five state-of-the-art gradient attacks: Fall of Empires (FOE) [43], A Little is Enough (ALIE) [5], Sign Flipping (SF) [2], Label Flipping (LF) [2], and Mimic [24].
In the Mimic attack, all Byzantine workers replicate a selected honest update and send it to the server. Inspired by the ALIE attack—which places Byzantine values between the honest mean and outliers to overcome robustness defenses—we instead select the honest outlier that is most surrounded by other honest updates (see Appendix for details). With this strategy, the “extreme” honest outliers hide Byzantine updates, which makes the attack harder to detect while remaining effective in heterogeneous settings. For instance, based on the extensive results in Appendix B.4, Tables 14, 15, and 16, Mimic often appears as the worst attack for the strategies –, BKT, and NNM.
All state-of-the-art strategies are evaluated using four robust aggregators: CWMed, CwTM, GM, and MKrum.
7.5 Metrics, Tables & Boxplots
We use two metrics: (i) Test Accuracy, averaged over training with a test set that matches the global training label distribution, and (ii) Gradient Dissimilarity, also averaged over training.
We ran experiments on four datasets, three levels of heterogeneity, four aggregation methods, and five attacks, averaging results over three random seeds for each scenario. Full results are in Appendix B.4. For clarity, we summarize key findings in three tables. Table 2 shows Test Accuracy, averaged over all aggregation methods (CWMed, CwTM, GM, MKrum) and all attacks (ALIE, FOE, LF, Mimic, SF), to reflect overall performance. Table 4 presents the same summary for Gradient Dissimilarity. Table 3 shows Test Accuracy under the strongest attack (i.e., the one causing the lowest performance), averaged over all aggregators, to highlight worst-case performance.
In Tables 2 and 3, we report the Accuracy Gain Over Top Non-WoLA Methods—the difference between each WoLA variant and the best-performing non-WoLA method (among –, BKT, FoundFL, and NNM). These values highlight performance gains over state-of-the-art strategies.
We also include boxplots of the gains from Tables 2 and 3 in Figures 4(a) and 4(b), to provide an overall view. In the following, when describing a given situation, we refer to the typical performance as the interquartile range of the corresponding boxplot. The average gain is also taken from these boxplots.
7.6 Optimization Algorithm
WoLA relies on approximating the global class gradients using local datasets (see Subsection 5.3). Approximation accuracy improves with increased data per class at each optimization step. As a result, the best performance is likely achieved with non-stochastic gradient descent, where honest workers compute gradients over the entire dataset. In practice, however, stochastic methods are more common—workers typically compute gradients on random mini-batches. To match this setting, we follow [3] and adopt the distributed stochastic heavy ball method (see Algorithm in Appendix B.1). This method has demonstrated strong robustness against Byzantine behavior in both heterogeneous [3] and homogeneous settings [13, 25, 16].
We use momentum $\beta=0.9$ for MNIST and Fashion-MNIST and $0.99$ for CIFAR10 and Purchase100.
7.7 Models & Hyperparameters
On MNIST and Fashion-MNIST, we train a lightweight CNN for $T=800$ steps. On CIFAR10, we use a deeper CNN; on Purchase100, a fully connected neural network. Both models are trained for $T=2000$ steps. The model architectures are provided in Appendix B.2. Unless otherwise specified, the loss is cross-entropy. The batch size is $b=128$, learning rates decay over time, gradient clipping is fixed at $5$, and $\ell_{2}$ regularization is set to $10^{-4}$.
7.8 Reproducibility
All experiments are run using the three seeds $\{1,2,3\}$ to ensure reproducibility. The code will be made publicly available for reuse at https://anonymous.4open.science/r/WoLA-E4CB.
8 Empirical Results
  • RQ1
    Does WoLA guide training toward a better model?
  • RQ2
    Are WoLA and WoLA more robust overall than state-of-the-art strategies?
  • RQ3
    Are WoLA and WoLA more robust in worst-case scenarios than state-of-the-art strategies?
  • RQ4
    Does combining WoLA or WoLA with NNM improve robustness?
  • RQ5
    What explains the robustness of WoLA methods?
  • RQ6
    How do batch size and momentum affect WoLA performance?
[…]
8.1 Targeted Model with WoLA (RQ1)
As expected from Proposition 3 (Subsection 6.2), WoLA does not target a better model during training.
8.2 Robustness of WoLA and WoLA (RQ2)
WoLA and WoLA significantly improve robustness. They perform slightly below the best state-of-the-art methods when $f=2$, offer marginal gains at $f=4$, and outperform all non-WoLA methods as the number of Byzantine workers increases.
When $f=2$: WoLA and WoLA show similar performance, slightly below the best non-WoLA strategies. Typical negative gains between $-1.1$ and 00 points (Figure 4(a)). Notable exceptions exist, such as Purchase100 with high heterogeneity (Table 2, $\alpha=0.3$), where NNM outperforms WoLA and WoLA by more than $6$ points.
When $f=4$: WoLA and WoLA still perform similarly and yield slightly better results than non-WoLA methods, with typical gains of a few points (Figure 4(a)). According to Table 2, gains range from $+5.4$ (CIFAR10, $\alpha=0.3$) to $-2.9$ (Purchase100, $\alpha=0.3$).
When $f=6$: WoLA slightly surpasses WoLA, and both outperform other methods. WoLA shows typical gains between $+3.3$ and $+12.1$ points, while WoLA reaches between $+1.8$ and $+8.1$ (Figure 4(a)). Table 2 reports maximum gains of $+21.4$ (WoLA) and $+20.5$ (WoLA) on MNIST with $\alpha=0.3$.
When $f=8$: WoLA is better than WoLA, and both methods clearly outperform all others (Figure 4(a)). WoLA typically achieves gains between $+5.6$ and $+14.7$ points (median around $+10.1$, Figure 4(a)), while WoLA ranges from $-3.7$ to $+7$ (median around $+5.2$, Figure 4(a)). The best results appear on MNIST with medium heterogeneity ($\alpha=1$): $+17.3$ for WoLA and $+13.6$ for WoLA (Table 2).
Importantly, when $f>2$, Table 2 shows that WoLA almost always ranks in the top three regardless of the scenario (exceptions: Purchase100 $\alpha=0.3$, $f=4$ and CIFAR10 $\alpha=0.3$, $f=8$). While WoLA does not frequently appear among the top values, it outperforms non-WoLA strategies most of the time on MNIST and Fashion MNIST, often on Purchase100, and occasionally on CIFAR10.
8.3 Worst-Case Robustness of WoLA and WoLA (RQ3)
Considering the strongest attack for each approach, WoLA and WoLA consistently enhance robustness, almost always outperforming other state-of-the-art methods, with gains reaching up to $+41.9$ points (Table 3, MNIST, $\alpha=1$, $f=8$, WoLA).
For $f=2$: WoLA and WoLA perform similarly, slightly surpass non-WoLA methods with typical gains of a few points (Figure 4(b)), and up to $+11.3$ points on CIFAR10, $\alpha=0.3$ (Table 3).
For $f=4$: WoLA and WoLA continue to perform similarly, outperforming all baselines with median gains around $+6.3$ and $4.6$ points respectively (Figure 4(b)). According to Table 3, WoLA’s gains range from $+0.1$ (MNIST, $\alpha=3$) to $+28.8$ (CIFAR10, $\alpha=0.3$), while WoLA ranges from 00 (MNIST, $\alpha=3$) to $+27$ (CIFAR10, $\alpha=0.3$).
For $f=6$ or $8$: WoLA outperforms WoLA, and both clearly surpass all other methods (Figure 4(b)). WoLA achieves median gains around $+13.4$ points, WoLA around $+12.4$ (Figure 4(b)). Table 3 shows that WoLA’s gains range from $-0.8$ (MNIST, $\alpha=3$, $f=6$) to $+41.9$ (MNIST, $\alpha=1$, $f=8$); WoLA ranges from $-3.4$ (CIFAR10, $\alpha=0.3$, $f=8$) to $+41.3$ (MNIST, $\alpha=1$, $f=8$).
8.4 Robustness Gains from Combining WoLA and WoLA with NNM (RQ4)
WoLA+NNM consistently achieves the top or top-three performance in Table 2 as soon as $f>2$, reflecting strong robustness across settings. Gains over the best non-WoLA methods can reach around $20$ points (Table 2). For example, on MNIST with $\alpha=0.3$ and $f=8$, the gain is $+26.6$; on Fashion MNIST with $\alpha=0.3$ and $f=6$, $+19.8$; on Purchase100 with $\alpha=1$ and $f=8$, $+25.2$; and on CIFAR10 with $\alpha=1$ and $f=6$, $+18.5$. The median gain is $+10.4$ when $f=6$ and $+17.8$ when $f=8$ (Figure 4(a)). Considering worst attacks, improvements are even larger, with median gains around $+20$ for both $f=6$ and $f=8$ (Figure 4(b)).
WoLA+NNM also regularly reaches top-three values in Tables 2 and 3, confirming its robustness. It performs similarly to WoLA+NNM when $f\leq 4$, and slightly worse when $f>4$, but still outperforms WoLA alone and all non-WoLA strategies (Figure 4(a), 4(b)). Importantly, it turns nearly all negative gains of WoLA alone into positive ones (Table 2). For instance, on CIFAR10 with $\alpha=3$ and $f=8$, the score improves from $-4.4$ to $+6.7$; with $\alpha=1$ and $f=8$, from $-3.6$ to $+8.3$; and with $\alpha=0.3$ and $f=2$, from $-0.5$ to $+2.2$. Overall, median gains over non-WoLA methods are around $+8$ when $f=6$ or $8$ (Figure 4(a)). Considering worst attacks, the median gain is around $+13$ for both $f=6$ and $f=8$ (Figure 4(b)).
8.5 Sources of Robustness in WoLA Methods (RQ5)
Gradient dissimilarity is a key indicator of system robustness [24, 3]. As expected from Proposition 2 (Subsection 6.2), WoLA methods significantly reduce gradient dissimilarity, thereby enhancing robustness against Byzantine attacks.
@@ 8.5 Sources of Robustness in WoLA Methods (RQ5)
8.6 Impact of Batch Size & Momentum on WoLA (RQ6)

### v3

@@ Introduction
(a) Boxplot of Global accuracy gains.
(b) Boxplot of Worst-case accuracy gains.
Figure 1: Accuracy gains over the best non-WoLA methods under low, medium, and high heterogeneity settings ($\alpha=3$, $1$, and $0.3$, respectively), based on the Global and Worst-case subtables reported in Tables 1 (Fashion MNIST & CIFAR10) and 4 (MNIST & Purchase100). Legend: $\bullet\!$ WoLA $\bullet\!$ WoLA+NNM $\bullet\!$ WoLA $\bullet\!$ WoLA+NNM
Experimental Setup
This section outlines our experimental setup, including the training objectives of WoLA, baseline methods, and robust aggregators. We describe the datasets and heterogeneity regimes, the distributed system configuration, Byzantine attack scenarios, evaluation metrics, optimization algorithms, models, and hyperparameters.
Training Objective, Baselines & Aggregators
We assume the goal is to learn a model that performs well on a test set matching the global label distribution. Accordingly, the ideal WoLA training objective aligns with this distribution, i.e., $q=p$.
We compare WoLA with four baselines: no pre-aggregation (-), Bucketing (BKT) (Karimireddy, He, and Jaggi 2020), Nearest-Neighbor Mixing (NNM) (Allouah et al. 2023), and FoundationFL (FoundFL) (Fang et al. 2025). We also evaluate WoLA combined with NNM. Methods requiring a server-side dataset or incurring substantial computational overhead(Karimireddy, He, and Jaggi 2020; Data and Diggavi 2021) are excluded.
To evaluate robustness to training objective attacks, we introduce WoLA, a worst-case scenario where Byzantine workers submit highly misleading label distributions. A simple robust aggregator is used to counter the attack (see Appendix, Section Training Objective Attack). We also assess WoLA combined with NNM.
All methods are tested with four robust aggregators: CWMed, CwTM, GM, and MKrum.
Datasets & Heterogeneity
We conduct experiments on four datasets: MNIST (LeCun et al. 1998), Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017), CIFAR10 (Krizhevsky 2009), and Purchase100 (Shokri et al. 2017). Due to space constraints, most results for MNIST and Purchase100 are provided in the Appendix.
To simulate data heterogeneity across workers, we sample from the global dataset using a Dirichlet distribution with concentration parameter $\alpha$, following the method in (Allouah et al. 2023; Hsu, Qi, and Brown 2019). We consider three heterogeneity levels—low, medium, and high—corresponding to $\alpha\in\{3,1,0.3\}$. Each worker holds a local dataset of fixed size, equal to the global dataset size divided by the number of honest workers.
Distributed system & Byzantine attacks
We consider a distributed system with $n=17$ workers, among which $f\in\{2,4,6,8\}$ are Byzantine. We evaluate five state-of-the-art gradient attacks: Fall of Empires (FOE) (Xie, Koyejo, and Gupta 2020), A Little is Enough (ALIE) (Baruch, Baruch, and Goldberg 2019), Sign Flipping (SF) (Allen-Zhu et al. 2020), Label Flipping (LF) (Allen-Zhu et al. 2020), and Mimic (Karimireddy, He, and Jaggi 2020).
In the Mimic attack, all Byzantine workers replicate a carefully selected honest update. Inspired by ALIE (Baruch, Baruch, and Goldberg 2019), we select the honest outlier most surrounded by other honest updates (see Appendix). This makes Byzantine updates harder to detect while remaining effective in heterogeneous settings. Tables 1416 in Appendix in Appendix show that Mimic is often the most harmful attack for baseline methods such as –, BKT, and NNM.
Metrics & Results
We evaluate performance using two metrics averaged over training: Test Accuracy and Gradient Dissimilarity.
Each experimental scenario—defined by dataset, heterogeneity level, defense method, and attack—is averaged over three runs with seeds $\{1,2,3\}$.
Comprehensive results are provided in Appendix and summarized in the following tables: Table 1 presents Test Accuracy averaged over training for Fashion-MNIST and CIFAR10. The global accuracy corresponds to the mean performance across all aggregators (CWMed, CwTM, GM, MKrum) and all attacks (ALIE, FOE, LF, Mimic, SF). The worst-case accuracy refers to the lowest accuracy across attacks, averaged over aggregators. Table 4 provides analogous results for MNIST and Purchase100.
Table 1 and 4 include Accuracy Gain Over Top Non-WoLA Methods, computed as the difference between each WoLA variant and the best-performing non-WoLA baseline. These gains are also visualized as boxplots in Figures 1(a) and 1(b).
Table 3 shows Gradient Dissimilarity averaged over training, showing the mean across all aggregators and attacks.
Due to space constraints, Table 3 and Table 4 are provided in the Appendix.
Optimization & Models
WoLA relies on approximating the global class-wise gradients with local class-wise gradients. The more data per class available at each optimization step, the better this approximation. As a result, the best performance is likely achieved using non-stochastic gradient descent, where honest workers compute gradients over the entire dataset. However, in practice, stochastic methods are more common. To align with this setting, we follow (Allouah et al. 2023) and adopt the distributed stochastic heavy ball method (see Algorithm in Appendix). This method has shown strong robustness against Byzantine behavior in both heterogeneous (Allouah et al. 2023) and homogeneous settings (El Mhamdi, Guerraoui, and Rouault 2021; Karimireddy, He, and Jaggi 2021; Farhadkhani et al. 2022). We use momentum $\beta=0.9$ for MNIST and Fashion-MNIST, and $\beta=0.99$ for CIFAR10 and Purchase100.
We use cross-entropy as point-wise loss, batch size $b=128$, learning rates that decay over time, gradient clipping at $5$, and $\ell_{2}$ regularization of $10^{-4}$.
We train a lightweight CNN for $T=800$ steps on MNIST and Fashion-MNIST. For CIFAR10, we use a deeper CNN, and for Purchase100, a fully connected network. Both are trained for $T=2000$ steps (details in Appendix).
Experimental Results
RQ 1 Are WoLA and WoLA more robust overall than state-of-the-art strategies?
RQ 2 Are WoLA and WoLA more robust in worst-case scenarios than state-of-the-art strategies?
RQ 3 Does combining WoLA or WoLA with NNM improve robustness?
In Appendix, we also answer to questions:
RQ A Does WoLA guide training toward a better model?
RQ B What explains the robustness of WoLA methods?
RQ C How do batch size and momentum affect WoLA performance?
Global Robustness of WoLA & WoLA (RQ 1)
According to Tables (a) and (b) in Table 1, WoLA and WoLA significantly improve robustness. The performance gains over the best state-of-the-art methods are particularly significant when the number of Byzantines is large.
When $f=2$: WoLA and WoLA show similar performance, slightly below the best non-WoLA strategies. Typical negative gains between 1.1-1.1 and 0 points (Figure 1(a)).
When $f=4$: WoLA and WoLA still perform similarly and yield slightly better results than non-WoLA methods, with typical gains of a few points (Figure 1(a)). According to Table 1, WoLA achieves gains of up to +5.4+5.4 (CIFAR10, $\alpha=0.3$).
When $f=6$: WoLA slightly surpasses WoLA, and both outperform other methods. WoLA shows typical gains between +3.3+3.3 and +12.1+12.1 points, while WoLA reaches between +1.8+1.8 and +8.1+8.1 (Figure 1(a)). Table 1 reports maximum gains of +16.6+16.6 (WoLA) and +10.3+10.3 (WoLA) on Fashion MNIST with $\alpha=0.3$.
When $f=8$: WoLA is better than WoLA, and both methods clearly outperform all others (Figure 1(a)). WoLA typically achieves gains between +5.6+5.6 and +14.7+14.7 points (median around +10.1+10.1, Figure 1(a)), while WoLA ranges from 3.7-3.7 to +7+7 (median around +5.2+5.2, Figure 1(a)). An exception occurs with CIFAR10, where both methods perform worse than NNM: 2.1-2.1 for WoLA and 6.2-6.2 for WoLA.
Importantly, when $f>2$, Table 1 shows that WoLA almost always ranks in the top three regardless of the scenario (exceptions: Fashion MNIST $\alpha=0.3$, $f=2$; and CIFAR10 $\alpha=0.3$, $f=2,8$). While WoLA rarely ranks among the top values, it usually outperforms non-WoLA strategies.
Worst-Case Robustness of WoLA & WoLA (RQ 2)
According to Tables (c) and (d) in Table 1, Considering the strongest attack for each approach, WoLA and WoLA consistently enhance robustness, almost always outperforming other state-of-the-art methods, with gains reaching up to +28.8+28.8 points (CIFAR10, $\alpha=0.3$, $f=4$, WoLA).
For $f=2$: WoLA and WoLA perform similarly, slightly surpass non-WoLA methods with typical gains of a few points (Figure 1(b)), and up to +11.3+11.3 points on CIFAR10, $\alpha=0.3$ (Table 1).
For $f=4$, WoLA and WoLA continue to perform similarly, both outperforming all baselines with median gains of approximately +6.3+6.3 and +4.6+4.6 points, respectively (Figure 1(b)). WoLA achieves gains of up to +28.8+28.8 (CIFAR-10, $\alpha=0.3$), while WoLA reaches up to +27.0+27.0 under the same setting.
For $f=6$ or 88: WoLA outperforms WoLA, and both clearly surpass all other methods (Figure 1(b)). WoLA achieves median gains around +13.4+13.4 points, WoLA around +12.4+12.4 (Figure 1(b)). Table 1 shows that WoLA achieves gains of up to +28.8+28.8, and WoLA up to +27.0+27.0 (CIFAR-10, $\alpha=0.3$, $f=4$).
Robustness Gains from Combining WoLA & WoLA with NNM (RQ 3)
WoLA+NNM consistently achieves the top or top-three performance in Table 1, reflecting strong robustness across settings. Gains over the best non-WoLA methods can reach around $20$ points. For example, on Fashion MNIST with $\alpha=0.3$ and $f=6$, +19.8+19.8; and on CIFAR10 with $\alpha=1$ and $f=6$, +18.5+18.5. The median gain is +10.4+10.4 when $f=6$ and +17.8+17.8 when $f=8$ (Figure 1(a)). Considering worst attacks, improvements are even larger, with median gains around +20+20 for both $f=6$ and $f=8$ (Figure 1(b)).
WoLA+NNM also regularly reaches top-three values in Tables 1, confirming its robustness. It performs similarly to WoLA+NNM when $f\leq 4$, and slightly worse when $f>4$, but still outperforms WoLA alone and all non-WoLA strategies (Figure 1(a)1(b)). Importantly, it turns nearly all negative gains of WoLA alone into positive ones (Table 1). For instance, on CIFAR10 with $\alpha=3$ and $f=8$, the score improves from 4.4-4.4 to +6.7+6.7; with $\alpha=1$ and $f=8$, from 3.6-3.6 to +8.3+8.3; and with $\alpha=0.3$ and $f=2$, from 0.5-0.5 to +2.2+$2.$ Overall, median gains over non-WoLA methods are around +8+8 when $f=6$ or $8$ (Figure 1(a)). Considering worst attacks, the median gain is around +13+13 for both $f=6$ and $f=8$ (Figure 1(b)).
Conclusion and Future Work
We proposed WoLA, a simple and effective loss reweighting method to improve robustness in federated learning under label skew. Unlike existing approaches that act after gradient computation, WoLA addresses the root cause of gradient divergence by aligning worker objectives through a shared training distribution. It can be combined with existing defenses like NNM for additional gains.
Empirically, WoLA significantly outperforms state-of-the-art methods: on CIFAR-10 with medium heterogeneity ($\alpha=1$, $f=6$), it improves accuracy by up to +11.3+11.3 points over the best non-WoLA baseline. On Fashion MNIST ($\alpha=0.3$, $f=6$), the gain is +16.6+16.6 points. Combined with NNM, WoLA+NNM achieves up to +18.5+18.5 on CIFAR-10 ($\alpha=1$, $f=6$), and +19.8+19.8 points more on Fashion MNIST ($\alpha=0.1$, $f=6$). Worst-case scenarios show even larger gains.
Theoretically, we prove that WoLA aligns honest gradients asymptotically, reducing dissimilarity to zero even under label skew. This holds regardless of the training objective.
Future work includes improving robustness further by refining the weighting scheme—e.g., at the mini-batch level—and designing stronger defenses against manipulations of the training objective.
Appendix A Notations
All notations used in the paper are summarized in Table 2.
Notation Meaning
$[k]$ Set of integers from $1$ to $k$, with $k\in\mathbb{N}$
#\# Cardinality of a set
$\mathbf{1}$ Indicator function
$\nabla$ Gradient
$\|\cdot\|$ $\ell_{2}$-norm
$\mathbb{E}[X]$ Expected value of a random variable $X$
$\mathbb{E}[X|\cdot]$ Conditional expectation of $X$
i.i.d. Independent and identically distributed
a.s. Probabilistic property known as almost surely
[…]
Notation Meaning
$\Phi$ Model
$n$ Number of workers
$\mathcal{H}$ Set of honest workers, with $\mathcal{H}\subset[n]$
$f$ Number of Byzantine workers, $f=n-\#\mathcal{H}$
Notation Meaning
$C$ Number of classes
$\mathcal{D}$ Global training set composed of observation-label pairs $(x,y)\in\mathcal{X}\times[C]$
NN Size of the global dataset: $N=\#\mathcal{D}$
$N_{i}$ Size of worker $i$’s local dataset: $N_{i}=\#\mathcal{D}_{i}$
Notation Meaning
$\mathcal{L}$ Loss
$\ell(y,\Phi(x))$ Pointwise loss of prediction $\Phi(x)$ and label $y$
$\mathcal{L}_{i}$ Worker $i$ loss: $\mathcal{L}_{i}=\sum_{(x,y)\in\mathcal{D}_{i}}\frac{\ell(y,\Phi(x))}{N_{i}}$, with $\sum_{i\in\mathcal{H}}\mathcal{L}_{i}\ N_{i}/N=\mathcal{L}$
$\mu^{c}$ Class-wise gradient of class $c$ in $\mathcal{D}$, i.e., $\sum_{(x,y)\in\mathcal{D}}\mathbf{1}(y=c)\nabla\ell(c,\Phi(x))/N^{c}$
$\mu^{c}_{i}$ Class-wise gradient of class $c$ in $\mathcal{D}_{i}$, i.e., $\sum_{(x,y)\in\mathcal{D}_{i}}\mathbf{1}(y=c)\nabla\ell(c,\Phi(x))/N^{c}_{i}$
$\mathcal{WL}$ Worker Label Alignement Loss (WoLA)
$\mathcal{W}\!\mathcal{L}_{i}$ Worker $i$ WoLA, $\sum_{i\in\mathcal{H}}\mathcal{W}\!\mathcal{L}_{i}\ N_{i}/N=\mathcal{WL}$
$q$ Shared training objective: a probability distribution $q=(q^{1},q^{2},\ldots,q^{C})$, i.e., $\sum_{i\in[C]}q^{i}=1$ and $q^{i}\geq 0$.
Appendix B Difference in Class-Wise Gradients
This section illustrates that class-wise gradients differ. Similar observations have been reported in previous studies (Zhang et al. 2022; Bao, Wu, and He 2024).
Figure 2 shows standard (non-FL) training of a shallow neural network on the Iris dataset (Fisher 1936). The loss decreases and accuracy improves over epochs, indicating successful model convergence. Independently of the optimization, at each step we compute the class-wise gradients Setosa, Versicolour, and Virginica, and track the cosine similarity between these gradients over training steps.
The figure shows that angles between class-wise gradients often exceed $\pi/2$, meaning the gradients point in very different directions. For Versicolour and Virginica, the gradients are nearly opposite. This highlights that different classes can push the model in distinct—and sometimes conflicting—directions during training.
Figure 2: Training a shallow neural network on the Iris dataset using batch gradient descent. The dataset is split into 70% training and 30% testing. While the model successfully learns Iris patterns, the class gradient directions vary significantly between classes. Legend: $\bullet\!$ Loss $\bullet\!$ Accuracy Gradient Cosine Similarity — $\bullet\!$ Setosa vs. Versicolour $\bullet\!$ Setosa vs. Virginica $\bullet\!$ Versicolour vs. Virginica
Appendix C Importance Sampling
The WoLA formula is derived from Importance Sampling, a Monte Carlo method that approximates an expectation under a target distribution by a weighted average of samples drawn from a different distribution (Tokdar and Kass 2010). In the following, we provide the detailed derivation of WoLA based on Importance Sampling.
Let $(X,Y)$ be random variables with joint density $f$, where realizations represent points in the global dataset $\mathcal{D}$. Let $q$ be the marginal distribution of $Y$. Similarly, let $(X_{i},Y_{i})$ be random variables with joint density $f_{i}$, representing points in the local dataset $\mathcal{D}_{i}$, and $p_{i}$ the marginal distribution of $Y_{i}$. The label distribution skewness assumption states that the conditional distributions of features given labels are identical across workers, i.e., $X|Y$ and $X_{i}|Y_{i}$ follow the same law.
From this, we have:
$\frac{f(x,y)}{\int f(x,y)dx}=\frac{f_{i}(x,y)}{\int f_{i}(x,y)dx}$
$\Leftrightarrow$ $\frac{f(x,y)}{q(y)}=\frac{f_{i}(x,y)}{p_{i}(y)}$
$\Leftrightarrow$ $\frac{f(x,y)}{f_{i}(x,y)}=\frac{q(y)}{p_{i}(y)},$
which leads to:
$\mathbb{E}\Big{(}\nabla\ell(Y,\Phi(X))\Big{)}$ $=\int\nabla\ell(y,\Phi(x))f(x,y)dxdy$
$=\int\nabla\ell(y,\Phi(x))\frac{f(x,y)}{f_{i}(x,y)}f_{i}(x,y)dxdy$
$=\int\nabla\ell(y,\Phi(x))\frac{q(y)}{p_{i}(y)}f_{i}(x,y)dxdy$
$=\mathbb{E}\Big{(}\nabla\ell(Y_{i},\Phi(X_{i}))\frac{q(Y_{i})}{p_{i}(Y_{i})}\Big{)}$
As a result, to estimate $\mathbb{E}\big{(}\nabla\ell(Y,\Phi(X))\big{)}$ using data drawn from the distribution of $(X_{i},Y_{i})$, we have to estimate $\mathbb{E}\big{(}\nabla\ell(Y_{i},\Phi(X_{i})){q(Y_{i})}/{p_{i}(Y_{i})}\big{)}$. This corresponds to the probabilistic form of the WoLA formula.
Appendix D Additional Research Questions
This section addresses additional research questions.
@@ Appendix D Additional Research Questions
Figure 3: Test Accuracy (averaged over three seeds) without Byzantine workers ($f=0$), using the mean aggregator on MNIST, Fashion MNIST, Purchase100, and CIFAR10, under low, medium, and high heterogeneity ($\alpha=3,1$, and $0.3$). Each column represents a dataset, and each row a heterogeneity level. Legend: $\bullet\!$ WoLA $\bullet\! -$
(a) Results under ALIE attack.
(b) Results under LF attack.
Figure 4: Test Accuracy and Gradient Dissimilarity (averaged over three seeds) under ALIE and LF attacks on Fashion MNIST with high heterogeneity ($\alpha=0.3$) and $6$ Byzantine workers ($f=6$). Each column corresponds to a robust aggregator, and each row to an evaluation metric. Legend: $\bullet\!$ WoLA $\bullet\!$ WoLA+NNM $\bullet\!$ WoLA $\bullet\!$ WoLA+NNM $\bullet\!$ NNM $\bullet\!$ FoundFL $\bullet\! -$ and BKT
@@
Targeted Model with WoLA (RQ A)
As expected from Proposition 3 in Subsection WoLA Properties, WoLA does not target a better model during training.
Sources of Robustness in WoLA Methods (RQ B)
Gradient dissimilarity is a key indicator of system robustness (Karimireddy, He, and Jaggi 2020; Allouah et al. 2023). As expected from Proposition 2 in Subsection WoLA Properties, WoLA methods significantly reduce gradient dissimilarity, thereby enhancing robustness against Byzantine attacks.
Impact of Batch Size & Momentum on WoLA (RQ C)